The Amount of H2O formed when 2g H2 and 1g O2 are reacted as, 2H2 + O2 → 2H2O ,will be 

A) 3 g
B) 1.125 g 
C) 4.5 g 
D) 2.50 g

Grade:11

3 Answers

Arun
6 years ago
Molar mass H2 = 2 g/mol 
Moles H2 = 2 g / 2 g/mol = 1
Molar mass O2 = 32 g/mol 
Moles O2 = 1 / 32 = 0.03125
2 H2 + O2 >> 2 H2O 
the ratio between H2 and O2 is 32 : 1 so we get 1 mole of water => 18 g
Arun
6 years ago
Sorry
For above answers' calculation,
Moles of H2 required = 0.03125 *2 = 0.06250
Hence moles of H2O formed = 0.06250
In gram = 0.0625 *18 = 1.125 gm
Option B is correct.
Regards
Arun (askIITians forum expert)
 
 
Jadhav
5 years ago
 
Sorry
For above answers' calculation,
Moles of hydrogen required = 0.03125 into 2 = 0.06250
Hence moles of water formed = 0.06250
In gram = 0.0625 *18 = 1.125 gm.